OpenStudy (anonymous):

MEDAL!!!!!!!!!!!!!! For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

OpenStudy (texaschic101):

Sally and Jenny decided to go to the store to buy some snacks. Sally purchased 2 candy bars and 1 bag of chips totaling $10. Jenny purchased 3 candy bars and 4 bags of chips totaling $25. How much did each candy bar (x) and each bag of chips (y) cost ? 2x + y = 10 ---> y = -2x + 10 now sub -2x + 10 in for y in the other equation 3x + 4y = 25 3x + 4(-2x + 10) = 25 3x - 8x + 40 = 25 3x - 8x = 25 - 40 - 5x = - 15 x = 3 now sub 3 in for x in either equation to find y 2x + y = 10 2(3) + y = 10 6 + y = 10 y = 10 - 6 y = 4 So each candy bar (x) costs $ 3 and each bag of chips (y) costs $4. Well...kind of expensive candy bars and chips....you can change them to something else if you want...thats up to you.

OpenStudy (anonymous):

thanks!

OpenStudy (texaschic101):

sure thing :)